if four hens lay 9 eggs in 7 days, how many days will it take four dozen hens to lay 27 dozen eggs?

Answer by stanbon(75887)

if four hens lay 9 eggs in 7 days, how many days will it take four dozen hens to lay 27 dozen eggs?
--------------------------------
hens and days are inversely related
days and eggs are directly related
------
d = k*e/h
----
Solve for "k" using "four hens lay 9 eggs in 7 days":
7 = k*9/4
---
k = 28/9
-----------
Equation:
d = (28/9)*e/h
------
how many days will it take four dozen hens to lay 27 dozen eggs?
d = (28/9)27/48
----
d = 28*3/48
---
days = 1 3/4 days
=======================
Cheers,
Stan H.